### v1

# ViGoR: Improving Visual Grounding of Large Vision Language Models with Fine-Grained Reward Modeling

Abstract
By combining natural language understanding and the generation capabilities and breadth of knowledge of large language models with image perception, recent large vision language models (LVLMs) have shown unprecedented reasoning capabilities in the real world. However, the generated text often suffers from inaccurate grounding in the visual input, resulting in errors such as hallucinating nonexistent scene elements, missing significant parts of the scene, and inferring incorrect attributes and relationships between objects. To address these issues, we introduce a novel framework, ViGoR (Visual Grounding Through Fine-Grained Reward Modeling) that utilizes fine-grained reward modeling to significantly enhance the visual grounding of LVLMs over pre-trained baselines. This improvement is efficiently achieved using much cheaper human evaluations instead of full supervisions, as well as automated methods. We show the effectiveness of our approach through numerous metrics on several benchmarks. Additionally, we construct a comprehensive and challenging dataset specifically designed to validate the visual grounding capabilities of LVLMs. Finally, we plan to release our human annotation comprising approximately 16,000 images and generated text pairs with fine-grained evaluations to contribute to related research in the community.
1 Introduction
Figure 1: An illustration of inaccurate visual grounding in Large Vision Language Models. BLIP-2 failed to provide detailed image descriptions while LLaVA offers more detail but lacks accuracy and frequently results in hallucinations. In contrast, our model preserves the logical reasoning and creativity of large language models while exhibiting significantly better accuracy and detail.
Large language models (LLMs) have garnered intense interest throughout the research and academic communities. Typically, these models are pre-trained with tremendous amounts of automatically aggregated text data and are further fine-tuned with specific user examples or evaluation feedback to enable the models to follow human-provided prompts and retrieve useful relevant information or solve logical problems. Recently, numerous techniques have augmented these breakthroughs with the ability to understand visual information by further integrating image features into the prompt encoding process. Although these works have successfully aligned image features into the large language model domain, they still exhibit severe problems. While a strong contextual grounding of language-only models can be learned from the enormous corpus of text data, training data for multimodal language/vision models is much more limited and flawed, while the complexity of the task is arguably higher. Large, automatically compiled datasets, such as the LAION-5B dataset [27], tend to feature only simple images with very short text descriptions. Training with such data often leads large vision language models to fail in capturing the essential details of the image, returning only a short and coarse description (see BLIP-2 [16] output in Figure 1). Moreover, techniques such as InstructBLIP [9] and BLIP [16] primarily rely on high-quality paired language/image datasets (e.g. VQAv2 [12], VizWiz [13], TextCaps [28], etc.). However, these datasets are expensive to collect and challenging to adapt for broader coverage due to the need for manual text annotation. On the other hand, initiatives such as LLaVA [20] train on perception and simple caption datasets in conjunction with the reasoning capabilities of LLMs to semi-automatically generate synthetic conversational ground truth. Unfortunately, such outputs can also contain nonfactual statements suffering from hallucinations, omissions, and inaccuracies in attribute or relational descriptions, as they are generated by text-only models based on sparse information about the actual image. Hence, the resulting trained model is still not ideal (see LLaVA’s output in Figure 1).
Instead, we propose a novel and generally applicable framework using fine-grained reward modeling. It efficiently and substantially enhances the visual grounding of LVLMs beyond pre-trained baselines such as LLaVA, while simultaneously preserving their capability to generate extended and detailed descriptions. Given a pre-trained LVLM (e.g., LLaVA), we input a set of images with prompts and sample multiple text outputs. Human annotators evaluate each image-text pair and assign fine-grained, per-sentence evaluation scores based on the quality of the generated text. This process results in a newly compiled dataset comprising image-text-evaluation trios. We train a reward model with this dataset and use it to fine-tune the pretrained LVLM (for further details, see Section  3). Notably, this fine-tuning process markedly improves the model’s visual grounding capabilities with just 16K data samples, demonstrating the method’s efficiency.
To further improve the performance of our system at negligible cost, we also develop an automatic method to build the reward model without additional human effort, which proves to be highly effective in improving the visual grounding capabilities of LVLMs.
Finally, we amalgamated the strengths of both reward models to develop a complete solution, which we refer to as ViGoR (Visual Grounding Through Fine-Grained Reward Modeling) throughout the remainder of the paper. To assess ViGoR’s efficacy, we evaluate it against two commonly used benchmarks, POPE [18] and MME [11], where it demonstrated significant advancements over baseline models. Moreover, we construct a comprehensive and challenging dataset, MMViG, specifically designed to validate the visual grounding capabilities of large vision language models. Finally, we use the recently released GPT-4V to automatically assist in evaluating our method.
In summary, our contributions are fourfold.
  • We introduce a novel framework that incorporates fine-grained reward modeling, substantially enhancing the visual grounding of LVLMs.
  • We develop reward models that require little human effort while leveraging the impressive advances in powerful and robust visual perception, demonstrating marked improvements in visual grounding efficiency.
  • We construct a comprehensive and challenging dataset, MMViG, specifically designed to validate the visual grounding capabilities of large vision language models.
  • We created and plan to release the human evaluation dataset comprising 16K pairs of images and generated results, along with fine-grained evaluation annotation providing detailed quality assessments.
2 Related Work
Large Vision Language Models
In the domain of large vision language models (LVLMs), recent advances have been remarkable. The integration of large language models (LLM) such as GPT [4, 24], PaLM [8], BLOOM [32], LLaMA [30], and Vicuna [7] has significantly advanced the field. Flamingo [1] and its open source counterpart OpenFlamingo [3], along with IDEFICS, have been pivotal in integrating LLMs with vision-language pretraining, employing innovative techniques like gated cross-attention dense blocks. PaLI’s [6] research on the scaling of vision and language (V&\&&L) components across various tasks has been instrumental. Additionally, PaLM-E’s [10] extension of LMM to the embodied domain, and BLIP-2’s [16] introduction of the Querying Transformer (Q-former) to align image and language encoders mark significant progress. InstructBLIP [9] further enhances this approach. Otter’s [15] focus on improving OpenFlamingo’s instruction-following capabilities and MiniGPT-4’s [35] recommendation for a single project layer to align visual and linguistic models demonstrate its efficiency and capability. mPLUG-Owl [33] introduces a novel approach by aligning visual characteristics before fine-tuning the language model using LoRA [14]. LLaVA directly injects visual tokens into a pre-trained LLM and finetunes the model with synthetic conversations generated by GPT-4 using metadata and short captions as input. While these techniques primarily focus on architectural innovation for aligning image features within the large language model domain or leveraging extensive image-text data for instruction tuning, our work introduces a distinct and novel general framework designed to enhance the visual grounding capability of any LVLM.
Visual Perception Models
Recent advances in visual perception models have demonstrated remarkable proficiency in handling a variety of tasks in open world scenes. In particular, CLIP [25] has exhibited a robust capability for zero-shot classification. By reformulating object detection as a grounding problem and incorporating additional grounding data, GLIP [17] has achieved semantic alignment at both the phrase and region levels, leading to impressive open-set detection performance. GroundingDINO [21] represents another significant step, grounding the state-of-the-art transformer-based object detector DINO [5] with language pretraining for open-set generalization. This model demonstrates a strong ability to discern whether elements are present in a scene and accurately detect object counts.
However, such advanced visual grounding abilities have not yet been fully realized in current LVLMs. To bridge this gap, we devised an automatic method for building a reward model using these vision perception models. Our approach aims to distill their sophisticated visual grounding capabilities directly into LVLMs.
Reward Modeling
Recent progress in training Large Language Models (LLMs) has increasingly emphasized the importance of reward modeling. This approach often incorporates human feedback and reinforcement learning optimization strategies, such as Proximal Policy Optimization (PPO). This approach is crucial in refining the accuracy and contextual relevance of model outputs. For example, Askell et al.  [2, 23] highlighted the potential of using human feedback in the training of general language assistants, emphasizing the importance of aligning model responses with human standards and values. Recently, LLaMA-2 [31] introduced a novel rejection sampling strategy within reward modeling, claiming that it improves the generation of contextually appropriate high-quality responses.
However, in the realm of Large Vision Language Models (LVLMs), the application of reward modeling remains underexplored, with most existing work focusing predominantly on instruction tuning [20, 35]. One exception is LLaVA-RLHF which adapts the Reinforcement Learning from Human Feedback (RLHF) from the text domain to the task of vision-language alignment, where human annotators are asked to compare two responses and pin-point the more hallucinated one [29]. Drawing inspiration from LLaMA-2, we propose the integration of reward modeling with a rejection sampling strategy into the LVLM training framework. Different from LLaVA-RLHF which uses a sparse reward model, our aim is to leverage fine-grained reward models to substantially improve the visual grounding capabilities of LVLMs, potentially leading to more accurate and contextually relevant output in vision-language tasks.
Figure 2: Overview of our model training framework. Starting with an input image and a prompt to generate a detailed caption, we sample a number of responses from the LVLM. These responses are passed through two fine-grained reward signal generation branches (a. by leveraging state-of-the-art generalizable perception models and b. by using an LVLM-based reward model trained using annotator feedback). Finally, we c. combine the fine-grained assessment signals from both sources into a single reward score, and select the best sampled description. Finally, we use heuristics and byproducts from the automated scoring system to further refine this sample, and use it for supervised fine-tuning of the LVLM.
3 ViGoR: Visual Grounding Improvement Framework
Our primary goal is to increase visual grounding and reduce hallucinations while maintaining the strong intuitive reasoning and creative thought process of pre-trained LLMs and LVLMs. While it has been shown that high quality human annotations for supervised fine-tuning is the most straightforward approach for significantly improving the performance of LVLMs, it is cost-prohibitive for many application scenarios. As such, we wish to efficiently leverage annotator time and the latest advances in direct visual perception models such as powerful open-set object detectors.
To achieve this, we construct a system to fine-tune a base LVLM with rejection sampling, similar to LLaMA2 [31], allowing the model to improve through using intelligent ranking of its own sampled output. This requires a robust and perceptive scoring system for the generated text. In our work, we use two complementary solutions in parallel. In particular, we train a reward model to incorporate human annotator assessments of text outputs from the LVLM, and provide positive and negative assessments of the LVLM during training time with unlabeled examples. As well, we leverage an open-set object detector and heuristics to verify the existence or absence in the image of the named noun entities extracted from the generated descriptions. Finally, we combine these signals into a single reward score, and use it to select the best description among the initial samples. This sample undergoes additional refinement and is used for the supervised fine-tuning of the LVLM. We refer the reader to Figure 2 for a visualization.
3.1 Reward Modeling via Fine-Grained Human Feedback
Human Preference Data Collection.
We design a system to incorporate human judgment and preference — considered the most reliable ground truth signal — into our model training. First, we select a pretrained LVLM’s checkpoint state, and create image / caption pairs from the model using a nonzero temperature to balance factual rigor with creativity. Unlike other approaches that ask annotators to provide relatively sparse judgments for the LLM/LVLM’s generated results, we ask crowd-workers to provide fine-grained feedback at the sentence level. For each sentence containing errors, the annotator selects the nature of the inaccuracy from a predefined list including object hallucinations, type of attribute error, incorrect relationships between multiple objects, error in the mentioned location or unreasonable conjectures about the image. Furthermore, the annotator provides a judgment about the creativity of the sentence. The creativity means if the sentence gives a thoughtful and reasonable interpretation or extrapolation of the image. Finally, the annotator provides a holistic assessment of the overall description’s level of detail and identifies the missing elements in the scene. These requirements encapsulate our overall goal: to enhance the LVLM’s visual grounding across the entire image while maintaining the insight and creativity inherent in the pre-trained language decoder.
Reward Model Training.
Using the collected annotations, we fine-tune a dedicated LVLM as the reward model on the annotations using instruction tuning to judge the base LVLM’s generated results during training time. The reward model is trained to output a string which encodes the various scores given the underlying image and the LVLM’s output as the input. Unlike existing work, we train the model to provide sentence-level human evaluations. To provide the necessary reference context for scoring each sentence, we prepend the sentence with all preceding generated text and explicitly ask the reward model to score the last sentence in the given passage. For example, a typical prompt is “Assess the accuracy of the last sentence in the following description of this image, and return a single number.” Due to the fine-grained feedback provided by the annotators, these prompts result in either a positive response (when no errors are found in the sentence), or a detailed negative response (where one of several error types are found in the sentence).
As we show in the ablation studies, this method significantly improves the reward model’s capabilities to guide the fine-tuning of the LVLM. Furthermore, we see that the reward model trained with fine-grained feedback can better understand the link between errors in the descriptions and the image, resulting in superior fine-tuning performance.
3.2 Reward Modeling with Automatic Methods
While the preference annotations directly encapsulate human preferences and are cheaper than supervised fine-tuning annotations, the human effort is still non-negligible. This can limit the scale of potential datasets and subsequently the visual discerning capabilities of the learned reward model due to overfitting. To further improve the cost efficiency of the overall system, we leverage the advances in discriminative vision perception models to automatically score the grounding and fidelity of text generated by a LVLM on large quantities of unlabeled images. However, these discriminative models generally have structured input (such as images and semantic classes) and structured outputs (such as bounding boxes). On the other hand, LVLMs operate with unconstrained and unstructured input and output (free-form text). This gap prevents directly providing the LVLM’s output to the discriminative models for scoring. As such, we carefully design the scoring system shown in part a of Figure 2.
Beginning with a caption generated by the LVLM from an image, we first identify the individual nouns mentioned using standard named noun entity recognition. Next, we employ an open-set object detector, prompted with these nouns and the input image, to verify the existence of the objects. A correct identification is rewarded with a positive score, while hallucinated objects incur a penalty. For this purpose, we use Grounding DINO [21] for its strong performance across a wide variety of image domains. However, we note that while these detection models are adept at identifying the existence of objects, their ability to detect other types of errors is limited (e.g., object attribution, object relationship…). This limitation underscores the continued necessity for human-preference-based reward modeling.
3.3 Reward Score and Rejection Sampling
As both our reward model-based and automated methods provide fine-grained reward scores, we must design a strategy to combine the scores into a single score to enable rejection sampling at the description level. Note that this process embodies the same ambiguity faced by annotators in existing work, where they are asked to analyze complete text output in detail and rank them based on holistic preference. However, our solution combines the detailed analysis in a more consistent fashion.
We iterate through the samples and find the best sample to use as ground truth for supervised fine-tuning. As our primary focus is to reduce generation of erroneous descriptions, we use the signals that indicate errors as the primary selector. For each sample, we aggregate all negative signals from the two streams of description evaluation by normalizing their values with their respective variances and linearly combine them. We select the sample with the best penalty score as the best candidate, and use the positive scores as a tiebreaker when necessary.
3.4 Refinement Module
After selecting the best candidate from the rejection sampling process, we design a simple refinement module to further polish this selected candidate description to use as a regression target.
For each sentence within the description, we assess the presence of noun phrases deemed nonexistent by the reward modeling module. Should a sentence contain any such noun phrases, we eliminate that sentence entirely from the description. For instance, consider the description:“The image features a black Dodge Charger parked on a brick road next to a sidewalk. A person is walking in front of the car.” If the reward modeling module identifies “person” as a non-existent noun phrase, the sentence “A person is walking in front of the car.” is removed.
This refinement process, as evidenced in Table 5, significantly improves the visual grounding capability of the model. It effectively reduces the hallucination problem by removing inaccurate elements, thereby leading to notable improvements in the model’s performance.
3.5 Model Training
The resulting description from the refinement module is used as ground truth in supervised fine-tuning using the standard autoregressive objective as in the original LLaVA [20]. As will be demonstrated by the ablation studies, two signal sources are complementary and provide better results than either one alone. This overall process is visualized in Figure 2.
4 Experiments and Results
In this section, we analyze the effectiveness of our approach in comparison with competitive baselines, and delve into the contributions from each component in our design. Finally, we provide visualizations to qualitatively compare the output of fine-tuned model with that of its initial state. In the following, “ViGoR-AT” refers to our approach utilizing only reward modeling with automatic methods. “ViGoR-RM” represents our method employing reward modeling with fine-grained human feedback. “ViGoR-All” and “ViGoR” denotes our method with the combination of the reward modeling with automatic methods and fine-grained human feedback.
4.1 Testing Framework
We use the recently proposed LLaVA [20] as the base model to demonstrate the effectiveness of our fine-tuning strategies without modifications to its architecture. In particular, we select the variant of the LLaVA model with the pre-trained and frozen ViT-L / 14 @ 224px CLIP image encoder [25] and the pre-trained Vicuna v1.3 [7] with 7B parameters as the language model. Our method is not specific to any particular LVLM testing architecture or configuration. We use this configuration for both our base LVLM model and our learned reward model.
With computation efficiency in mind, we select the smallest variant of the Grounding DINO model with the Swin-T [22] backbone with the official checkpoint and the default box and text thresholds for the open-set object detection task. Our proposed method can likely directly benefit from the larger (but more computationally expensive) variants or future advances in these models.
4.2 LVLM Model Fine-tuning
To support our experiment, we used 25,574 images from the ADE20K dataset [34] (training split) as the basis, as they cover a wide variety of relatively complex scenes. To preserve the generalization of our method, we disregard the original annotations of the dataset. We elicit responses from our base LVLM through a set of question prompts (e.g. “Please provide a detailed description of the given image.”), and sample five different outputs for each input image. Subsequently, these outputs are fed into our reward model to obtain a corresponding reward score for each description. In addition to keeping the highest-scoring description following the common practice in rejection sampling, we enhance this description further by applying a refinement module to achieve additional quality improvements.
To reduce computational complexity, we choose an offline supervision generation strategy where we save the refined output as the ground truth for its respective image using the initial model state. However, we note that the rejection sampling process can take place online by refining an evolving model state. We believe that this has the potential to further improve the algorithm’s effectiveness, but requires significantly more computation.
During the fine-tuning phase, we set the learning rate at $2\times 10^{-5}$ and train the model over two epochs with a batch size of 32. The entire process is executed on eight 40G A100 GPUs, taking 7 hours in total to complete.
Method HL CA AA RA RL RS DL Average
LLaVA 3.28 3.27 3.28 3.27 3.27 3.27 3.01 3.24
ViGoR-AT 2.26 2.27 2.25 2.26 2.25 2.30 2.06 2.24
ViGoR-RM 2.48 2.50 2.47 2.49 2.49 2.46 2.96 2.55
ViGoR-All 1.98 1.96 2.00 1.98 1.99 1.97 1.97 1.98
Table 1: Detailed caption analysis via GPT-4Vision ranking. Scores are the average preference rank of the caption when compared within the 4 candidates in the table (ranging from 1 to 4; lower is better).
HL: Does the description have hallucinations?
CA: Does the description have good counting accuracy for objects?
AA: Does the description assign accurate attributes to objects?
RA: Does the description have accurate relationships between objects?
RL: Is the description relevant to the image as a whole?
RS: Does the description exhibit reasonable thought and reasoning?
DL: Does the description encompass all details in the image?
4.3 Reward Model Training
Human feedback collection.
To allow the reward model to be highly receptive to errors in the descriptions, we use a pre-trained checkpoint for LLaVA that is already fine-tuned with our automatic reward generation mechanism. Through hands-on experimentation, we observed that while the images from the ADE20K dataset exhibit excellent scene variability, the resulting cognitive load for human annotators is very high. As such, instead, we generate 16,000 detailed image captions using images selected from the somewhat simpler MS COCO [19] dataset. We enlist the services of a team of 15 professional annotation workers to assist us in the creation of the evaluation data set for the training of the reward model using the process described in Section 3.1. We provide carefully designed and detailed annotation instructions (available in our Supplementary Materials) to our annotators, along with extensive sample annotations. The process took approximately 3 weeks.
Reward model.
We initialize the same model architecture as our base LVLM with the same weights used to generate the annotation samples as the starting state of our reward model. We train this model for 5 epochs on the dataset with 16,000 samples with a batch size of 32 and an initial learning rate of $2\times 10^{-5}$.
Method #P VQA Random Popular Adversarial
AC PR AC PR AC PR
mPLUG-Owl 7B 53.3 51.7 50.6 50.3 50.7 50.3
MiniGPT-4 7B 77.8 75.4 68.3 64.3 66.6 62.45
MM-GPT 9B 50.0 50.0 50.0 50.0 50.0 50.0
InstructBLIP 13B 88.7 85.0 81.4 75.1 74.4 67.7
LLaVA 7B 54.4 52.3 52.4 51.3 50.8 50.4
ViGoR (ours) 7B 85.1 89.0 81.5 83.0 75.5 73.8
Improvement +30.7 +36.7 +29.1 +31.7 +24.7 +23.4
Table 2: Quantitative results on POPE benchmark.
4.4 Quantitative Results
The types of natural interactions that a user can have with a LVLM are essentially unlimited in variety. However, since the focus of our work is on improving visual grounding, we primarily evaluate the benefits of our technique in two types of settings. First, we analyze the unstructured output of the LVLM when it is asked to generate a detailed and comprehensive caption for an image. This tests the model’s ability to be perceptive to all regions and objects in the image. Second, we evaluate the model’s performance on responding to targeted questions for which a concise and unambiguous answer is required. This further tests the model’s ability to allow the input prompt to guide its visual understanding and formulate an accurate response.
Method #Param Data VQA Data EX CT PO CO PT CE SC LM AT OC Overall
MiniGPT-4 7B 5K 68.3 55.0 43.3 75.0 41.8 54.4 71.8 54.0 60.5 57.5 581.7
VPGTrans 7B 1.4M 70.0 85.0 63.3 73.3 84.0 53.5 141.8 64.8 77.3 77.5 790.5
InstructBLIP 13B 1.2M 185.0 143.3 66.7 153.3 123.8 101.2 153.0 79.8 134.3 72.5 1212.8
Cheetor 7B 500K 180.0 96.7 80.0 116.7 147.3 164.1 156.0 145.7 113.5 100.0 1299.9
Muffin 13B 267K 195.0 163.3 66.7 165.0 137.8 81.8 151.3 146.3 116.5 57.5 1281.0
LLaVA [^1] We report the number of LLaVA v1.0 reproduced by ourselves. 7B 158K 158.3 83.3 51.7 85.0 94.2 85.0 145.8 125.8 74.3 57.5 960.1
ViGoR (ours) 7B 174K 180.0 143.3 83.3 100.0 128.2 121.2 150.5 130.8 127.0 145.0 1309.3
Improvement +21.7 +60.0 +31.6 +15.0 +34.0 +36.2 +4.7 +5.0 +52.7 +87.5 +349.2
Table 3: Quantitative results on MME benchmark. “VQA Data” indicates whether the model is fine-tuned using the VQA datasets. For a fair comparison, we compare the Large Vision Language Model using the Vicuna backbone.
Method HL CA CO MA Overall
InstructBLIP 187.98 110.20 143.42 104.14 545.74
LLaVA 156.19 61.44 97.84 62.26 377.73
ViGoR-AT 183.81 106.34 124.36 122.08 536.59
ViGoR-RM 178.92 94.89 133.33 137.82 544.96
ViGoR-All 187.61 112.08 135.22 140.40 575.31
Table 4: Quantitative results on our MMViG benchmark.
4.4.1 GPT4-Assistant Evaluation
Objective quantitative measurement of free-form responses has been a major challenge in the LLM and LVLM space due to the inherent ambiguities in defining quality. Furthermore, it is unclear how structured information can be extracted from generated text so that a score can be assigned using traditional objective functions. To address this, we utilize the advanced capabilities of the GPT-4Vision service to evaluate the outputs. We ask GPT-4Vision to rank the output of different LVLM candidates according to numerous criteria that signify the quality of the captions. This offers a detailed view of the visual foundation and text generation capabilities of the LVLMs under test. We perform this evaluation with 96 randomly selected images from the COCO validation dataset and show the results in Table 1 along with an overview of the definitions of the metrics. We select the four most relevant configurations for comparison: 1) the original LLaVA baseline model that we use as the initialization; 2) ViGoR-AT: the ViGoR method using only reward modeling with automatic methods; 3) ViGoR-RM: the ViGoR method using only the guidance from the reward model trained using the annotator feedback, and lastly 3) ViGoR-All: the complete model. As is clearly evident, our complete model consistently achieves the best ranking within the comparison across all metrics. Furthermore, we observe that either source of supervision signals significantly outperforms the original baseline model. Further details are provided in the supplementary materials.
Figure 3: Qualitative results. We show examples of descriptions generated by our technique after using the fine-tuning scheme, and compare them with the output from the original LLaVA [20]. Our approach is able to greatly reduce the amount of hallucinations and invalid observations while increasing the amount of detailed visual descriptions. Furthermore, the model retains the plausible intuitive reasoning capabilities of the underlying LLM.
4.4.2 Short Answer Benchmark Evaluation
Numerous benchmarks have been released recently to support the development of LVLMs. To decrease the complexity and ambiguity, these benchmarks typically consist of questions with concise answers such as “Is there a red apple in the image?” As this style of questioning is not explicitly considered in our model training, these benchmarks measure both the visual grounding capabilities and the generalization of our method. As such, we select the commonly used POPE [18] and MME [11] benchmarks, which are constructed using images from the validation set of the COCO dataset [19]. POPE probes the LVLM with 3000 yes/no questions targeting 500 images to evaluate the LVLM’s ability to determine the existence of specific objects in a scene. MME has a similar questioning setup, but extends the questioning subject to cover aspects such as numerical count, position, and color of objects, as well as other vision tasks such as scene/landmark identification and OCR.
As illustrated in Tables 2 and  3, our model shows consistent improvement across all categories evaluated in both datasets. Furthermore, while the competitive baseline methods use VQA datasets containing ground truth for these types of straightforward questions, our method is able to achieve comparable or superior results without fine-tuning on such resources. This suggests that the visual grounding abilities learned by the model through our process is general. When the model learns to generate accurate comprehensive descriptions for images, it is able to apply the capabilities to related reasoning tasks, suggesting that the logical deduction capabilities of the underlying pre-trained LLM is retained and leveraged. Furthermore, our 7 billion parameter model shows on par performance with some models with nearly double the number of parameters (13 billion). This underscores the efficiency and robustness of our model in handling diverse visual-language tasks.
However, POPE only measures object-level hallucinations or omissions, while MME provides as few as 60 question/answer ground truth pairs over 30 images for the various categories. To increase the comprehensiveness of our analysis, we introduce a new benchmark, MMViG, to assess the visual grounding capabilities of LVLMs in four distinct categories: Hallucinations (HL), Counting Accuracy (CA), Color Accuracy (CO), Material Accuracy (MA). Our evaluation process involves selecting 200 images from the MS COCO 2017 validation data set for each category. We create a set of 400 paired yes/no questions across the four categories. We automatically generate the questions for HL, CA, CO and MA using ground-truth labels from both the COCO and PACO [26] annotations. This methodology allows us to thoroughly examine the LVLM’s ability to accurately interpret and respond to a diverse range of visual inputs. As shown in Table 4, the improvements and trends observed in previous scenarios are also evident in these assessments.
4.5 Qualitative Results
In Figure 3, we show qualitative results of our technique in action, focusing on a comparative analysis between our model and its starting point (LLaVA). This comparison is particularly relevant as our model undergoes fine-tuning based on the LLaVA model. By juxtaposing these two models, we highlight the enhancements and advancements our model has achieved, offering valuable insights into the efficacy and impact of our training strategies.
AT-RM HF-RM Refine MMViG
- - 521.41
- 536.59
- 544.96
- 551.87
575.31
Table 5: Ablation studies on different components of our framework.
4.6 Ablation Studies
In this section, we present ablation studies of our approach to justify various design choices. For simplicity, we evaluated all variants on the MMViG benchmark.
4.6.1 Ablation Studies of Framework Components
reward model design MMViG
none 377.73
holistic-based 403.01
fine-grained-based 544.96
Table 6: Fine-grained vs holistic evaluation by reward model.
In this section, we examine the efficacy of different components within our framework. Table 5 presents the performance of each module, underscoring their individual significance. The results clearly demonstrate that integrating all components yields the most optimal performance.
4.6.2 Fine-grained vs Holistic Evaluation by Reward Model
In this section, we perform a comparative assessment between the fine-grained analysis of sentences by the reward model as compared with the more commonly used holistic evaluation of complete responses. This assessment is carried out in the context of reward modeling with human feedback. For the holistic reward model, we retrained a reward model using a singular, unified reward for each description. As shown in Table 6, the fine-grain-based approach demonstrates better performance than the holistic approach, which shows that denser and more informative signals allow the reward model to better discern the quality of the generated text by creating more direct links between the visual features in the image and the sentences describing them.
5 Conclusion and Future Work
Our work takes a step toward improving visual grounding capabilities of LVLMs to reduce hallucination, errors in relational reasoning, counting, and so on. To this end, we developed a framework named ViGoR to combine fine-grained reward modeling of human preferences with powerful existing open-set visual perception models to efficiently improve LVLMs in these aspects. To validate our approach, we collect the first large dataset with fine-grained human annotation feedback for the LLaVA. We show that ViGoR significantly improves the LLaVA model’s ability to generate accurate and relevant text given input images, while preserving its ability to creatively and intuitively reason about the scene.
In the future, we plan to apply RLHF training with our ViGoR framework, which may offer additional improvements over our training based on rejection sampling. As well, we anticipate that explicit visual alignment predictions through reasoning about object and element localization can further improve the visual grounding of LVLMs.